# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v4

@@ III-C Crosshair detection and masking
((b)) Final crosshair detection after Hough line averaging.
Figure 3: Crosshair detection process.
If the crosshair is not detected or if there is a large deviation from the expected crosshair position, we use an average of the previous crosshair positions as estimate. In principle, our system can work without the laser diode by calibration one initial crosshair and using a virtual crosshair afterwards, which stays in the same position in the image under ideal conditions. Since we estimate the heading using only the $x$-direction chessboard lines and we need no crosshair line directions, this would mean fixating one point in the image and using it as AMR position. In the experiments with our low-cost robot that has limited mechanical stability, we have found that using the laser diode, which is lightweight and mounted much closer to the ground than the camera, is more accurate and robust than using a virtual crosshair due to less vibrations and shaking of the laser diode.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing, as shown in Fig. 4(b). The binary Canny output is then processed by Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation, which consists of erosion followed by dilation. In erosion and dilation, a 2D kernel is slided over the image, where the kernel center corresponds to the pixel to process. In each kernel, erosion only keeps white (detected line) pixels if all kernel pixels are white. Dilation does the opposite and sets the color of the center pixel to white if any pixel in the kernel is white. The image after Hough transformation and opening is depicted in Fig. 4(c). For further operations, the image is inverted, as seen in Fig. 4(d). The chessboard squares are detected by first finding and simplifying the white square contours with the functions findContours [31] and approxPolyDP [28, 9]. The simplified contours (polygons) are then filtered for square detection. Only polygons that have four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. As the detected squares are slightly smaller due to the line broadening and opening, 10 pixels are added to the square corner coordinates. Those detected squares are given in Fig. 4(e). In the last step, the detected corners are refined to subpixel accuracy with cornerSubPix [13] as shown in Fig. 4(f).
If squares are occluded, the redundancy in the chessboard pattern enables estimating all squares (including the square with crosshair used for positioning) by simply extending the detected squares to build virtual, undetected squares. Given the precise square detection, this method has proven quite accurate. If for some reason no square is detected, we apply a constant velocity model for $x$, $y$ and $\theta$ to estimate the AMR position and heading, which is sufficient for a high enough sampling frequency of camera images.
@@
((b)) Canny edges with masked crosshair.
((c)) Hough transform and opening.
((d)) Inversion.
((e)) Square detection (colored).
((f)) Corner refinement.
Figure 4: Square detection process in the perspective transformed top-down image.
@@ III-E Position and heading estimation
If no QR code is detected, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of four ambiguous angles (due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. The four ambiguous chessboard line directions are computed by clustering and averaging the lines from Hough transformation depicted in Fig. 4(c). Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this nearest neighbor assignment works reliably.
@@
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square while crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the unambiguous region (e.g. by increasing the square size), increase the sampling rate of the camera, or add a robust and accurate crosshair position prediction model (e.g. through odometry).
If no QR code is detected in the first frames, we assume an initial position and heading of $x,y,\theta]^{T}=[0,0,0]^{T}$ and then correct the initial estimates through rotation and translation once a QR code is available.
@@ IV-A Hardware
[…]
@@ IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since hand-collecting ground truth on the floor [3] (e.g. with a set square and a protractor) is an error prone and tedious task that only works in a static setting, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand. We therefore zoom in and select the crosshair center and three points of the respective square to build the coordinate system for positioning, including the $x$-direction line for heading estimation. The points are selected to the best of our abilities with pixel accuracy, where one pixel equals roughly $\frac{16.66}{300}\simeq 0.56$ mm (see Fig. 1(b)).
@@ V Results
The labeled and estimated trajectories are shown in Fig. 5. As visible, the estimates closely follow the ground truth.
Figure 5: Trajectory evaluation
@@
((b)) Cumulative density functions of the absolute errors.
@@
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). A few outliers arise due to small errors in the crosshair or corner detection, often caused by overlaps of the crosshair with the chessboard lines and corners. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
@@
$\theta$ -0.05 0.42 1.09 1.43 1.47
All estimates have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors and the max errors in $x$ and $y$ are well below 1 cm.
Heading estimation also shows good performance with the 95th percentile of absolute heading errors just slightly being over $1^{\circ}$ and a non-negligible part of this error likely being due to the labeling process. Good heading accuracy is important to accurately map from the crosshair center to another position around the crosshair, e.g. to a localization sensor on the AMR.
@@ VI Discussion
Cost and portability: Our system is low-cost due to the low hardware requirements (camera, chessboard, processor, optional laser diode). We are confident that the system will also work with even cheaper hardware. The largest cost factor is currently the chessboard floor. However, any grid pattern is sufficient. For rooms without such chessboard floor, one might e.g. build a grid by projection orthogonal laser lines on the floor using a laser level and drawing or taping such line pattern on the floor using the laser lines. One may also build a portable version, where pre-built chessboard grid patches can be sticked together. The system can further work with a cheaper camera and since the image processing is relatively simple, one might implement a future (online) version that runs on a field programmable gate array (FPGA).
Accuracy: Due to the high spatial resolution of modern cameras and the closeness of the camera to the position references (chessboard corners) and the crosshair on the ground, we achieve high positioning accuracy. The localization accuracy of GG is also mostly independent of the AMR position in the room, given similar illumination conditions. Performance might be enhanced by multi-camera and multi-laser setups. The localization method could be extended to handle arbitrary detected chessboard corners, opposed to the four square corners in which the crosshair lies. Our current positioning bottleneck (with a moving AMR) is the not the localization algorithm, but the mechanical stability of the AMR, which causes larger shaking and vibration of the camera, and to a smaller extend also of the laser diode and the crosshair. The resulting position and heading jitter is not considered in our evaluation. Simple hardware improvements include an AMR with better stability, more rigid camera and laser diode installations, and stabilizers for the camera and also for the laser diode in case of localization using a real crosshair. On the software side, smoothing the trajectory, e.g. online with an alpha-beta or a Kalman filter [2], or offline with a Savitzky-Golay [29] filter, would reduce this jitter.
@@
3D position and orientation estimation: An alternative localization approach that could be studied in future work is Perspective-n-Point (PnP)-based localization with a monocular camera, as often used with fiducial markers and realized in a conceptionally similar system to ours with a hand-held camera and a small, printed chessboard in [30]. Therefore, the known correspondences between the detected 2D chessboard corner points in the image and their known 3D positions in world coordinates enable 3D camera position and orientation estimation without the laser diode. While potentially less accurate in 2D position estimation, this could minimize the error from projecting the estimated position by GG to the position of another localization sensor on the AMR. The current localization method could theoretically employ a (virtual) crosshair located on the floor below the AMR for positioning. However, this requires installing the camera and additional lighting at the bottom of the AMR chassis and dealing with very small camera-ground distance (motion blur & visibility of full chessboard squares).
@@
Global localization: By equipping the floor with QR codes, we achieve absolute, global localization capabilities. QR codes were, however, just a readily available solution for this proof-of-concept of the system. Simpler visual encodings that are easier and faster to detect would improve the system, especially for an online version. The encodings are used to identify the chessboard square corners, e.g. with the known QR corners as in our current version, or by other means such as making the different grid lines identifiable.
Real-time processing: In addition to the implemented crosshair tracking and the proposed FPGA version, simpler global position and orientation references etc., the image processing still has much potential for improvement. Instead of running the full QR code, crosshair and corner detection for each image as described in Sec. III (which itself has room for improvement, e.g. by applying a simpler method run at lower resolution to find the initial chessboard corners instead of the expensive Hough transformation at full resolution), one might track the global references and the periodic chessboard corners after initial detection and make use of the prior knowledge about the crosshair position in the image. This would require processing only smaller patches of the images, thus speeding up the detection chain.
@@
Robustness: The redundant, periodic chessboard pattern allows our system to work even if large areas of the chessboard are occluded, e.g. by other AMRs, people, dirt etc. Since the crosshair stays roughly in the same position of the image and parts of the crosshair are sufficient for crosshair detection, we can also deal with crosshair occlusion. Robustness could be further increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy, but also increase system complexity and cost. Moreover, the crosshair could be replaced by a more redundant laser pattern.
Scalability: Our GG system is scalable to large areas. In addition, the system can be scaled to robot swarms. In a version without the laser diode (virtual crosshair or PnP), there will be no interference between robots except for some floor occlusion, which the system can handle. If all robots use the same laser diode, one can still differentiate the different laser crosshairs, as the crosshair of each respective robot stays in the same place in the images. In case there might be interference between the crosshairs, one could try installing lasers of different colors or patterns on the robots.
@@ VII Conclusion
We have introduced a novel, camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera placed at low height on the AMR, and a laser diode that projects a crosshair from the AMR onto the chessboard floor. The crosshair on the floor is continuously observed by the camera for position and heading estimation. Our prototype achieved low mm localization and sub-degree heading accuracy. We have further discussed the many advantages this localization approach brings and how it may be improved and extended in various directions in future work.

### v5

@@ III-C Crosshair detection
((b)) Final crosshair detection.
Figure 3: Crosshair detection.
If the crosshair is not detected or at an unexpected position, we use an average of the previous crosshair positions as estimate. In principle, our prototype can work without the laser diode by replacing the real crosshair with a virtual one, which stays in the same position in the image under ideal conditions (i.e. with a perfectly stabilized camera). This means selecting one point in image coordinates to replace the crosshair center, and using it as AMR position. Heading is estimated based on the chessboard rather than the crosshair orientation. On our low-cost robot, the lower mounted laser diode proved to be more accurate and robust than a virtual crosshair, being less affected by vibrations under AMR motion.
@@ III-D Chessboard square detection
To detect the chessboard squares, we first apply a Canny filter for edge detection after grayscale conversion and Gaussian smoothing. The binary Canny output is then processed by probabilistic Hough transformation to find the chessboard lines. For further denoising, the detected lines are broadened and the image of lines is processed by an opening operation. The image after Hough transformation and opening is depicted in Fig. 4(b). We then invert the image. The chessboard squares are detected by first finding and simplifying the white square contours. The simplified contours (polygons) are then filtered for square detection, where only polygons with four edges, relative side lengths between factor 0.95 and 1.05, and an area of 80.000 - 100.000 pixels are kept. Finally, the detected corners are refined to sub-pixel accuracy [6], shown in Fig. 4(c).
If squares are occluded, the redundancy in the chessboard pattern enables estimating all squares by simply extending the detected squares to build virtual, undetected squares. Using the precise square detection, this method has proven quite accurate. If for some reason no square is detected, we apply a constant velocity model for $x$, $y$ and $\theta$, which is sufficient with a high enough sampling frequency of camera images.
@@
((b)) Hough lines after opening.
((c)) Square detection (red).
Figure 4: Square detection in the top-down image.
@@ III-E Position and heading estimation
If no QR code is detected, we predict the current heading using a constant velocity model. The heading is then estimated by choosing the detected angle from the set of four ambiguous angles (due to the lack of knowledge which chessboard lines point in $x$- and $y$-direction) by a nearest neighbor assignment. The four ambiguous chessboard line directions are computed by clustering and averaging the lines from Hough transformation in Fig. 4(b). Since the AMR has to move in a physically plausible way and the heading cannot make large jumps between two images, this nearest neighbor assignment works reliably.
@@
Using this crosshair tracking after an initial global position and heading estimate with a QR code, further frequent global updates with the QR code references are not necessary, but this requires a high enough image sampling frequency such that the Nyquist-Shannon sampling theorem holds. The spatial frequency $f_{pattern}$ of the chessboard pattern is $f_{pattern}=\frac{1}{16.\overline{6}}\,\mathrm{cm}^{-1}$ and thus the images have to be sampled with a spatial frequency of at least $f_{image}=2f_{pattern}$ to avoid aliasing (causing jumps into the wrong square during crosshair tracking). This upper bounds the maximum AMR velocity to $v_{max}=f_{image}^{-1}f_{fps}$, where $f_{fps}$ is the sampling rate of the camera. We record with $f_{fps}=8.\overline{3}$ Hz, i.e. $v_{max}\approx 0.69\,\frac{\mathrm{m}}{\mathrm{s}}$, which is more than double of the maximum velocity of our AMR. To adapt the crosshair tracking for faster AMRs, one would have to either increase the unambiguous region (e.g. by increasing the square size), increase the sampling rate of the camera, or integrate an accurate kinematic model (e.g. through odometry).
If no QR code is detected in the first frames, we assume an initial position and heading of $x,y,\theta]^{T}=[0,0,0]^{T}$ and then correct the initial estimates through rotation and translation once a QR code becomes available.
@@ IV-A Hardware
Our prototype AMR is a low-cost Raspberry Pi robot (SunFounder Picar-X, $\sim\mathdollar 100$) with Ackermann kinematics. The USB RGB fisheye camera is from Svpro $(\sim\mathdollar 100)$ and offers a resolution of 1920x1200 pixels, a FoV of $150^{\circ}$, and fast image creation (i.e. negligible motion blur) by employing a global shutter. The green laser diode is from CTRICALVER $(<\mathdollar 10)$ with a wavelength of $520\text{\,}\mathrm{n}\mathrm{m}$ and has $1\text{\,}\mathrm{m}\mathrm{W}$ of power. The chessboard floor is made from Polyvinylchlorid (PVC) with a known square size of ($\overline{6}, 16.6¯1$$\overline{6}$) cm.
@@ IV-B Localization experiments
To showcase the correct working of our system under real inference conditions with possible motion blur, AMR vibrations etc., we drive the AMR in the laboratory by steering it wirelessly with a PS4 controller. Since we have no access to a costly high-accuracy reference system, which was our motivation to design the low-cost and high-accuracy GG, we developed a software tool to label the ground truth position and heading based on the corresponding top-down image by hand, similar to the annotation procedure of various localization approaches based on pixel differences [23, 20]. We therefore zoom into the wanted points and select the crosshair center position, and three corners of the respective square to build the coordinate system for localization, including the $x$-direction line for heading estimation. The points are selected to the best of our abilities with pixel accuracy, where one pixel equals roughly $\frac{16.66}{300}\approx 0.56$ mm (see Fig. 1(b)). Future work should benchmark GG against external ground truth, potentially in a version more robust to AMR vibrations discussed in Sec. VI.
@@ V Results
The labeled and estimated trajectories are shown in Fig. 5. The estimates closely follow our hand-labeled trajectory.
Figure 5: Trajectory evaluation.
@@
((b)) Cumulative distribution functions of the absolute errors.
@@
A closer error evaluation is found in Fig. 6. The errors in $x$, $y$ and $\theta$ are depicted in Fig. 6(a). A few outliers arise due to small errors in the crosshair or corner detection, often caused by overlaps of the crosshair with the lines and corners of the chessboard. The corresponding CDFs of the absolute errors are found in Fig. 6(b). They show clearly that most errors in $x$ and $y$ are in low mm or even sub-mm range.
@@
$\theta$ -0.05 0.36 0.87 1.30 1.39
All estimates have negligible bias. Positioning errors in both dimensions are similar as expected. The errors in $x$ are higher due to more outliers as visible in Fig. 6(a). The 99th percentile of the absolute errors and the max errors in $x$ and $y$ are well below 1 cm. Heading estimation also shows good performance with the 95th percentile of absolute heading errors being slightly under $1^{\circ}$. Good heading accuracy is important to accurately map from the crosshair center to another position around the crosshair, e.g. to a localization sensor on the AMR.
@@ VI Discussion
Cost and portability: The largest cost factor of GG is the chessboard floor, which is still cheaper than using other mm-accuracy methods for small rooms. Rooms without a chessboard floor could instead use a taped or drawn grid of lines. By connecting pre-built chessboard patches, one could also port the system.
Accuracy: Positioning accuracy is mostly independent of the AMR position in the room, given similar illumination conditions. Current localization requires four crosshair square corners and might be extended to accept any (and less) corners. This allows crosshair projection closer to the AMR, or replacing the crosshair center with a known, visible point on the AMR (see e.g. Fig. 2(a) bottom) for position referencing. The error in mapping from GG’s position to a sensor on the AMR could also be minimized with a camera mount beneath the AMR. However, this reduces FoV and the feasible chessboard square size. It also increases motion blur, necessitating reduced exposure time and thus additional lighting or camera flash. The current limitation during motion is the mechanical stability of the AMR, causing camera and crosshair jitter. Since we only evaluate the crosshair position accuracy, not a jittered projection to a stable position on the AMR, this effect is currently not measured. Such jitter could be reduced with more stable AMRs, rigid camera and laser diode mounts, mechanical stabilizers, and by smoothing the trajectory. Replacing the crosshair center with a position reference on the AMR, or with a stabilized point in image coordinates, makes the laser diode obsolete.
@@
3D positioning: Future work could study Perspective-n-Point (PnP)-based localization, as often used with fiducial markers. The known camera intrinsics and correspondences between the detected 2D image chessboard corner points and their known 3D world positions enable 3D camera position and orientation estimation. PnP requires no laser diode and might also reduce the error from projecting the estimated position by GG to the position of another localization sensor on the AMR.
@@
Global localization: By placing QR codes on the floor, GG achieves absolute, global localization. QR codes were, however, just a convenient solution for this proof-of-concept. Simpler visual encodings with faster detection would improve the system, especially for online operation. The encodings serve to identify the chessboard square corners, e.g. with the known QR corners as in our current version, or by alternatives such as identifiable grid lines.
Real-time processing: Instead of running the full detection chain for each image as described in Sec. III, one could track the global references and chessboard corners after initial detection. This requires processing only smaller image patches, thus speeding up the detection chain.
@@
Robustness: The redundant, periodic chessboard pattern allows GG to operate even when large areas of the chessboard are occluded, e.g. by other AMRs, people, dirt etc. Robustness could be increased by adding more cameras, laser diodes, and odometry, which would additionally improve accuracy, but also increase system complexity and cost.
@@ VII Conclusion
We have introduced a camera-based indoor localization system called GroundGazer (GG) for 2D AMR position and heading estimation. Opposed to currently existing methods with similar mm localization and sub-degree heading accuracy, our system has much lower hardware requirements. Its limitation is the need for a chessboard or a similar grid pattern and global position and orientation references on the floor, which may limit its usability for applications where this is impractical. We built a prototype of GG using a chessboard floor with known corner coordinates as passive position references, QR codes as global position and orientation references, a ground gazing fisheye camera placed at low height on the AMR, and a laser diode that projects a crosshair from the AMR onto the chessboard floor. The crosshair and the floor are continuously observed by the camera for position and heading estimation. Our prototype achieved low mm localization and sub-degree heading accuracy. We have further discussed the advantages this localization approach brings and how it may be improved and extended in various directions in future work.